OpenStudy (anonymous):

ok im stuck on this one 1000+m^3

OpenStudy (jdoe0001):

so, what are you supposed to do?

OpenStudy (anonymous):

factor it

OpenStudy (jdoe0001):

ok, recall that \(\bf a^3+b^3 = (a+b)(a^2-ab+b^2)\)

OpenStudy (jdoe0001):

now notice the 1000, does it look like the cube of something?

OpenStudy (anonymous):

yes isnt 1000 a and m^3 b

OpenStudy (anonymous):

no it dose not look like a cube or anything

OpenStudy (jdoe0001):

well, that's was just an equation to factor about :/

OpenStudy (jdoe0001):

notice the 1000, does it look like the cube of some number?

OpenStudy (jdoe0001):

could be write 1000 as say \(\bf \square?^3\) ?

OpenStudy (anonymous):

i dont see a cube or anything for the 1000 do i have to find numbers that = to 1000

OpenStudy (jdoe0001):

ok,, how much is 10 x 10 x 10?

OpenStudy (anonymous):

ok thats 1000 so i change the 1000 to 10

OpenStudy (jdoe0001):

\(\bf 1000+m^3 \implies 10^3+m^3\\ a^3+b^3 = (a+b)(a^2-ab+b^2)\)

OpenStudy (anonymous):

ok so it would be 10^3+m^3=(10+m) (10^2-2m+m^2)

OpenStudy (jdoe0001):

-ab that is -(10)(m) \(\bf \ne\) -2

OpenStudy (jdoe0001):

but otherwise, yes

OpenStudy (jdoe0001):

-(10)(m) \(\bf \ne \) -2m that is